Two charges (+3µC on the left of the -5µC) are separated by a distance of 5um on a horizontal line. Point P is located 10µm above the 3µC charge.


a) Draw a sketch of the setting discussed in the question above. Make sure to label all points discussed. (2marks)


b) Calculate the field at point P (3marks)


c) Calculate the potential at point P 

(1 mark)


1
Expert's answer
2021-10-18T11:13:48-0400

(a) Let "q_1=+3\ \mu C", "q_2 = -5\ \mu C". Let the distance between charges "q_1" and "q_2" be "r_{12}=5\ \mu m", the distance from the charge "q_1" to the point P above it be "r_{1P}=10\ \mu m". Let's draw a sketch:



(b) The electric field due to point charge "q_1" at point P directed upward (away from the charge "q_1"), the electric field due to point charge "q_2" at point P directed toward charge "q_2".

Let’s first find the magnitudes of the electric fields "E_1" and "E_2":


"E_1=\dfrac{k|q_1|}{r_{1P}^2}=\dfrac{9.0\cdot10^{9}\ \dfrac{N\cdot m^2}{C^2}\cdot|3\cdot10^{-6}\ C|}{(10\cdot10^{-6}\ m)^2}=2.7\cdot10^{14}\ \dfrac{N}{C},""E_2=\dfrac{k|q_2|}{r_{2P}^2}=\dfrac{9.0\cdot10^{9}\ \dfrac{N\cdot m^2}{C^2}\cdot|-5\cdot10^{-6}\ C|}{\sqrt{(10\cdot10^{-6}\ m)^2+(5\cdot10^{-6}\ m)^2}}=4.0\cdot10^{9}\ \dfrac{N}{C}."

Let’s find projections of the fields "E_1" and "E_2" on axis "x"- and "y":


"E_{1x}=0, E_{1y}=E_1=2.7\cdot10^{14}\ \dfrac{N}{C},""E_{2x}=E_2cos\theta=E_2\dfrac{r_{1P}}{r_{2P}},""E_{2x}=4.0\cdot10^{9}\ \dfrac{N}{C}\cdot\dfrac{10\cdot10^{-6}\ m}{\sqrt{(10\cdot10^{-6}\ m)^2+(5\cdot10^{-6}\ m)^2}}=3.6\cdot10^9\ \dfrac{N}{C},""E_{2y}=E_2sin\theta=E_2\dfrac{r_{12}}{r_{2P}},""E_{2y}=4.0\cdot10^{9}\ \dfrac{N}{C}\cdot\dfrac{5\cdot10^{-6}\ m}{\sqrt{(10\cdot10^{-6}\ m)^2+(5\cdot10^{-6}\ m)^2}}=1.78\cdot10^9\ \dfrac{N}{C}."

Then, we can find "x"- and "y"-components of the net electric field at point P:


"E_x=E_{2x}=3.6\cdot10^9\ \dfrac{N}{C},""E_y=E_{1y}-E_{2y}=2.7\cdot10^{14}\ \dfrac{N}{C}-1.78\cdot10^9\ \dfrac{N}{C}=2.69\cdot10^{14}\ \dfrac{N}{C}."

We can find the net electric field at point P from the Pythagorean theorem:

"E_{net}=\sqrt{E_x^2+E_y^2}=\sqrt{(3.6\cdot10^9\ \dfrac{N}{C})^2+(2.69\cdot10^{14}\ \dfrac{N}{C})^2}=2.69\cdot10^{14}\ \dfrac{N}{C}."

We can find the direction of the net electric field from the geometry:


"\theta=tan^{-1}(\dfrac{E_y}{E_x})=tan^{-1}(\dfrac{2.69\cdot10^{14}\ \dfrac{N}{C}}{3.6\cdot10^9\ \dfrac{N}{C}})=89.9^{\circ}."

(c) We can find the potential at point P as follows:


"V=V_1+V_2=\dfrac{kq_1}{r_{1P}}+\dfrac{kq_2}{r_{2P}},""V=\dfrac{9.0\cdot10^{9}\ \dfrac{N\cdot m^2}{C^2}\cdot3\cdot10^{-6}\ C}{10\cdot10^{-6}\ m}+\dfrac{9.0\cdot10^{9}\ \dfrac{N\cdot m^2}{C^2}\cdot(-5\cdot10^{-6}\ C)}{\sqrt{(10\cdot10^{-6}\ m)^2+(5\cdot10^{-6}\ m)^2}},""V=-1.32\cdot10^{9}\ V."